Problem 11-1A (Video) Partially correct answer. Your answer is partially correct. Try again. Rogen Corporation manufactures a single product. The standard cost per unit of product is shown below. Direct materials—1 pound plastic at $6.00 per pound $ 6.00 Direct labor—2.0 hours at $11.90 per hour 23.80 Variable manufacturing overhead 12.00 Fixed manufacturing overhead 20.00 Total standard cost per unit $61.80 The predetermined manufacturing overhead rate is $16.00 per direct labor hour ($32.00 ÷ 2.0). It was computed from a master manufacturing overhead budget based on normal production of 10,400 direct labor hours (5,200 units) for the month. The master budget showed total variable costs of $62,400 ($6.00 per hour) and total fixed overhead costs of $104,000 ($10.00 per hour). Actual costs for October in producing 4,100 units were as follows. Direct materials (4,240 pounds) $ 26,288 Direct labor (8,030 hours) 97,163 Variable overhead 99,710 Fixed overhead 33,390 Total manufacturing costs $256,551 The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored. (a) Compute all of the materials and labor variances. Total materials variance $Entry field with correct answer -1688 Entry field with correct answer Materials price variance $Entry field with correct answer -848 Entry field with correct answer Materials quantity variance $Entry field with correct answer -840 Entry field with correct answer Total labor variance $Entry field with correct answer 417 Entry field with correct answer Labor price variance $Entry field with correct answer -1606 Entry field with correct answer Labor quantity variance $Entry field with correct answer 2023 Entry field with correct answer (b) Compute the total overhead variance. Total overhead variance $Entry field with incorrect answer 100300 Entry field with correct answer Click if you would like to Show Work for this question: Open Show Work
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
4240 $6.20 $26,288.00 4240 $6.00 $25,440.00 4100 $6.00 $24,600.00
$848.00 U $840.00 U
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $848.00 U
Direct material quantity variance $840.00 U
Direct material cost variance $1,688.00 U
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
8030 $12.10 $97,163.00 8030 $11.90 $95,557.00 8200 $11.90 $97,580.00
$1,606.00 U $2,023.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $1,606.00 U
Direct Labor Efficiency variance $2,023.00 F
Direct labor cost variance $417.00 F

Solution b:

Total overhead variance = Overhead applied - Actual overhead

= (8200*$16) - ($99,710 + $33,390)

= $1,900 U